OpenStudy (anonymous):

{-n^2} from n=0 to inf , according to text this sequence is bounded above and the upper bound value is 0, now should the upper bound value must be between the given limit that's for this question is n=0 to inf therefore the upper bound value has come to zero if not then the upper bound value could be anything so why not we say it could be -1,-2,-3,-4,...,-inf ?

OpenStudy (anonymous):

an upper bound is zero an upper bound is also 22 or any positive number

OpenStudy (anonymous):

how come? -n^2 > -n^2 +1

OpenStudy (anonymous):

you wrote \[\{-n^2\}_{n=0}^{\infty}\] so the seqence should look like \[\{-1,-4,-9,-16,...\}\]

OpenStudy (anonymous):

exactlly

OpenStudy (anonymous):

\[\{0,-1,-4,-9,-16,...\}\] sorry that was wrong, it should start at n = 0

OpenStudy (anonymous):

ok SO the upper or lower bound value should be from the sequence?

OpenStudy (anonymous):

no not necessarily

OpenStudy (anonymous):

i just made a mistake, since it says start at n = 0 i need to replace n by 0 first

OpenStudy (anonymous):

I mean if the upper and lower limit has been defined then it should be from the sequence? isn't?

OpenStudy (anonymous):

so 0 IS an upper bound, but the upper bound is not unique the least upper bound is uniqe, and that is 0

OpenStudy (anonymous):

there is clearly no lower bound for this one, since any number you pick you can always find a number \[-n^2\] less that it

OpenStudy (anonymous):

do we plugin the bound value in sequence to test if it works?

OpenStudy (anonymous):

if "a" is an upper bound it should mean \[a\geq x_i\] for all \[x_i\in \{x_n\}\]

OpenStudy (anonymous):

mostly you find a upper bound by thinking

OpenStudy (anonymous):

thinking what? :|

OpenStudy (anonymous):

not by a computation, although you may use one if you have a sequence \[\{0,-1,-4,-9,-16...\}\] it is pretty clear that 0 is an upper bound, and in fact that 0 is the least upper bound

OpenStudy (anonymous):

if you have say \[\{\frac{1}{n}\}_{n=1}^{\infty}\] you can see that the least upper bound is 1, and the greatest lower bound is 0

OpenStudy (anonymous):

do we have to think that if we plugin the bound value in the sequence formula and it satisfies the condition of lower or upper bound values

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how did you find the bound value so quickly? is it because it is n=1 to inf?

OpenStudy (anonymous):

if you want an upper bound you are trying to find an "a" greater than or equal to all numbers in the sequence

OpenStudy (anonymous):

sure just by thinking

OpenStudy (anonymous):

{(-1)^n+1} from n 1 to inf the upper bound value is 1 and lower bound value is -1 how did we found these values?

OpenStudy (anonymous):

sequence starts \[1,\frac{1}{2},\frac{1}{3},...\] clearly it goes to zero and clearly 1 is the largest

OpenStudy (anonymous):

okkkkkkkkkkkkkkkkkkk

OpenStudy (anonymous):

your sequence takes on only two values, 1 and -1

OpenStudy (anonymous):

so yes, 1 is the largest and -1 is the smallest

OpenStudy (anonymous):

if a equals to every number of the sequence then will it still be the upper bound value?

OpenStudy (anonymous):

again it comes from looking at the terms in the sequence and see what you have

OpenStudy (anonymous):

think of the silly sequence \[\{1,1,1,1,...\}\] not a very interesting sequence, a constant sequence. the upper bound is 1, and likewise the lower bound is 1

OpenStudy (anonymous):

so all we have to analyse the sequence values?

OpenStudy (anonymous):

why the upper and lower are both same in this silly seq?

OpenStudy (anonymous):

because for every element x in the squence both of these statements are true: \[1\leq x\] and \[1\geq x\]

OpenStudy (anonymous):

ok the theoram says it not only have to be greater or less it could be equal as well

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanksssss bro

OpenStudy (anonymous):

yw